-2x+1=3(x+2)
Solve for x and simplify. HELP this is the last one.

Answer by texttutoring(324)

-2x+1=3(x+2)
Get rid of the brackets on the right-hand side by multiplying each term by 3.
-2x+1=3x+6
Add 2x to both sides:
-2x+1+2x=3x+6+2x
1=5x+6
Subtract 6 from both sides:
1-6=5x+0
-5=5x
Divide both sides by 5:
x=-1